OpenStudy (anonymous):

Find the inverse Laplace transform of F(S)= (2e^(-3s))/(s(s-1)(s-2))

OpenStudy (anonymous):

\[ F(S)= \frac{ 2e^{-3s} }{s(s-1)(s-2) }\]

OpenStudy (anonymous):

First thing I would take care of is the exponential factor using the formula, \[\mathcal{L}\{u(t-c)f(t-c)\}=e^{-cs}\mathcal{L}\{f(t)\}\]

OpenStudy (anonymous):

Applying this, you have \[\mathcal{L}^{-1}\left\{\frac{2e^{-3s}}{s(s-1)(s-2)}\right\}=u(t-3)f(t-3)\] where \(f(t)=\mathcal{L}^{-1}\{F^*(s)\}\), with \(F^*(s)=\dfrac{1}{s(s-1)(s-2)}\). As for finding \(f(t)\), break up into partial fractions.

OpenStudy (anonymous):

@SithsAndGiggles I am confused by your last part. How did you get the right side? \[\mathcal{L}^{-1}\left\{\frac{2e^{-3s}}{s(s-1)(s-2)}\right\}=u(t-3)f(t-3) \]

OpenStudy (anonymous):

\[\mathcal{L}\{u(t-c)f(t-c)\}=e^{-cs}\mathcal{L}\{f(t)\}\] implies \[u(t-c)f(t-c)=\mathcal{L}^{-1}\left\{e^{-cs}\mathcal{L}\{f(t)\}\right\}\] In this case, you have \(c=3\), and \[\color{red}{e^{-cs}}\color{blue}{\mathcal{L}\{f(t)\}}=\frac{\color{blue}{2}\color{red}{e^{-3s}}}{\color{blue}{s(s-1)(s-2)}}\]

OpenStudy (anonymous):

Makes much sense now. Thanks